OpenStudy (anonymous):

What is square root 27x^3? Assume x represents a positive value

OpenStudy (anonymous):

(3*x)^1.5

OpenStudy (anonymous):

\[\sqrt{27x^3}\]

OpenStudy (anonymous):

\[\sqrt{27x^3}=\sqrt{9\times 3\times x^2\times x}=3x\sqrt{3x}\]

OpenStudy (anonymous):

\[3\sqrt(3)\sqrt(x^3)=3\sqrt(3x^3)\]